OpenStudy (canada907cat):

Find the value of d^2y/(dx^2) for the hyperbola defined by the equation y^2-x^2=7 at the point (3,4). I know that the first derivative equation is x/y and is equals 3/4 at the point of (3,4) but I do not know how to find the second derivative.

OpenStudy (eliesaab):

Find the second derivative of \[ f(x)=\sqrt{x^2+7} \]

OpenStudy (eliesaab):

\[ f'(x)=\frac{x}{\sqrt{x^2+7}} \]

OpenStudy (eliesaab):

\[ f''(x)=\frac{7}{\left(x^2+7\right)^{3/2}} \]

OpenStudy (eliesaab):

\[ f''(3)=\frac 7{64} \]

OpenStudy (eliesaab):

You can also do it by implicit differentiation

OpenStudy (eliesaab):

First you find \[ y'(x)=\frac{x}{y(x)} \]

OpenStudy (eliesaab):

then you find \[ y''(x)=\frac{y(x)-x y'(x)}{y(x)^2} \]

OpenStudy (eliesaab):

Replace x by 3, y'(x) by 3/4, y(x) by 4 you obtain the same result

OpenStudy (canada907cat):

Okay so would that look like 4-3(3/4)/16^2? @eliesaab